# Assessing treatment efficacy for interval-censored endpoints using multistate semi-Markov models fit to multiple data streams

### v1

@@ 5 Discussion
Our computational framework is closest in spirit to the Bayesian data augmentation MCMC algorithm of Barone and Tancredi (2022). Like us, they use a Markov surrogate process to propose sample paths for a semi-Markov process conditionally on the data. Bayesian inference is attractive in settings where multistate models are applied because we often have information about aspects of disease progression that can be leveraged to resolve aspects of the model that are not strongly identifiable in the data. However, the task of incorporating prior information about observables is not easily translated into priors about model parameters for complicated models. Bayesian data augmentation MCMC also presents with computational challenges as convergence and mixing deteriorate when the data become sparse Papaspiliopoulos and others (2007). In these settings, a computationally expensive, albeit often robust, solution is pseudo-marginal MCMC (Andrieu and Roberts, 2009), which integrates over the missing data within the MCMC. An attractive alternative that could sidestep the awkwardness of prior elicitation and thorny computational issues is to employ an optimization-centric approach to Bayesian inference, such as the weighted Bayesian bootstrap (Newton and others, 2021). Our computational framework would naturally fold into such an approach and would turn the problem of posterior sampling into a set of optimization problems that can be solved in parallel while allowing for prior information about observables to be encoded via penalized likelihoods.
@@
Multistate semi-Markov models encode more realistic biological assumptions than Markov models that have historically been applied in analyzing data under panel observation. Our computational framework is general, rather than model specific, and allows for misclassification in the observed state. This innovation is critical to our analysis of the REGEN-2069 data. We employ importance sampling to recycle sample paths within an MCEM algorithm, thus minimizing computational overhead. We further explore semi-parametric inference via splines, but leave the door open for further work on penalized splines. It would be interesting to explore likelihoods that jointly penalize the flexibility of all semi-parametric components, as in Machado and others (2021).
@@
The multistate models explored in this work were formulated under the assumption that the observation process is independent of the underlying disease process, which is reasonable in REGEN-2069 trial. However, it is commonly recognized that failing to accommodate for non-ignorability of the observation process, as would arise from informative sampling times or disease driven observation (DDO) processes, can confound inference about the disease process (Lange and others, 2018; Cook and Lawless, 2021). It would be straightforward to extend our framework to accommodate DDO processes by modifying the likelihood and proposal algorithm to account for preferential sampling given covariates or the latent disease state. The proposal algorithm could also be easily modified to accommodate more complicated observation schemes, such as when data indicate that a state was not visited over some time interval, e.g., in the ACTT-1 study of remdesivir for treatment of COVID-19, the data recorded the highest-intensity level of respiratory therapy over the preceding 24 hours. This could be addressed by setting intensities to zero for transitions into states that are ruled out by the data, thus assuring that impossible trajectories would not be considered. The MCEM algorithm described in this paper can also be extended by substituting the Markov proposal model with a richer proposal model from which it is straightforward to sample paths conditional on the panel data. Candidates include discrete mixtures of Markov models (Elvira and others, 2019) and phase-type continuous-time Markov chains (Titman and Sharples, 2010). These richer proposal would have the potential of making the importance sampling estimate of the marginal log-likelihood, and thereby the MCEM algorithm as a whole, more robust.
@@ Acknowledgments
The authors would like to thank to Michael P. Fay, Jing Wang, Arthur Winer, and Jennifer Gillman for their invaluable help that greatly improving this manuscript. This work utilized the computational resources of the NIH HPC Biowulf cluster (https://hpc.nih.gov). This project has been funded in whole or in part with federal funds from the National Cancer Institute, National Institutes of Health, under Contract No. 75N91019D00024. The content of this publication does not necessarily reflect the views or policies of the Department of Health and Human Services, nor does mention of trade names, commercial products, or organizations imply endorsement by the U.S. Government. Conflict of Interest: None declared.

### v2

@@ 5 Discussion
Multistate semi-Markov models encode more realistic biological assumptions than Markov models that have historically been applied in analyzing data under panel observation. Our computational framework is general, rather than model specific, and allows for misclassification in the observed state. This innovation is critical to our analysis of the REGEN-2069 data. We employ importance sampling to recycle sample paths within an MCEM algorithm, thus minimizing computational overhead. We further explore semi-parametric inference via splines and leave the door open for further work on penalized splines. It would be interesting to explore likelihoods that jointly penalize the flexibility of all semi-parametric components, as in Machado and others (2021). We also regard optimization-centric Bayesian approaches, such as the weighted Bayesian bootstrap (Newton and others, 2021), to inference for multistate models in biomedical applications as an important avenue for future work since there is a need to use prior information to resolve aspects of the model that are not identified in the data. Our framework would naturally fold into such an approach, which turns the problem of posterior sampling into a set of optimization problems that can be solved in parallel while allowing for prior information about observables to be encoded via penalized likelihoods or pseudo-observations.
Our computational framework is closest, in spirit, to the Bayesian data augmentation MCMC algorithm of Barone and Tancredi (2022), who also employ a Markov surrogate to propose semi-Markov sample paths conditionally on the data. While data augmentation MCMC may be tractable for fitting simple models to small datasets, it is difficult to see this as a general solution in more complicated modeling scenarios when it becomes increasingly expensive to sample paths at each MCMC iteration. Furthermore, convergence and mixing deteriorate as the data become increasingly sparse relative to dynamics of the process (Papaspiliopoulos and others, 2007). These problems are compounded when poor convergence and mixing necessitate longer MCMC runs to obtain an adequate posterior sample. Nevertheless, we caution that our MCEM algorithm is not a panacea for non-identifiability despite its being agnostic about the structure of the model and accommodating a variety of measurement processes. General conditions under which MLEs for intermittently observed Markov processes may fail to exist, or are practically not identifiable, are characterized in Bladt and Sørensen (2005). The intuition from these results, such as needing sufficient temporal granularity in the data to ensure practical identifiability, applies equally to semi-Markov models. Scenarios that are computationally challenging for maximum likelihood estimation often prove difficult for Bayesian computation in the absence of well-calibrated priors. In our experience, our algorithm tends to fail “loudly” in settings where a responsible data analyst would reasonably expect that a model is over-parameterized given the extent of the data. We view obvious computational failure as a strength of the algorithm since it indicates that the model is not supported by the data.
We have limited our exposition to multistate models that are progressive and where the observation process is independent of the underlying disease process, both of which are reasonable in the context the REGEN-2069 trial. It is possible, in theory, to apply our MCEM algorithm to models with back-transitions. In practice, fitting models for recurrent processes under intermittent observation is difficult and often requires knowledge about the number of back-transitions to be brought to bear by restricting the state space of the model or by appealing to informative priors or a penalized likelihood. Thus, we defer exploration of recurrent models for future work. We also highlight extensions towards automatic smoothing parameter selection in penalized maximum likelihood estimation as being of particular importance given the need to mitigate the influence of modeling choices that could have undue influence on model fit, e.g., knot placement for regression splines. Penalized maximum likelihood estimation was explored in the Markov case in Machado and others (2021) and it would be of great interest to adapt aspects of their approach to fitting semi-Markov models via MCEM.
@@
It is also commonly recognized that failing to accommodate for non-ignorability of the observation process, as would arise under a disease driven observation (DDO) processes, can confound inference (Lange and others, 2018; Cook and Lawless, 2021). It would be straightforward to extend our framework to accommodate DDO processes by modifying the likelihood and proposal algorithm to account for preferential sampling. The proposal algorithm could also be easily modified to accommodate more complicated observation schemes, such as when data indicate that a state was not visited over some time interval, e.g., in the ACTT-1 study of remdesivir for treatment of COVID-19 the data recorded the highest-intensity level of respiratory therapy over the preceding 24 hours. This could be addressed by setting intensities to zero for transitions into states that are ruled out by the data, thus assuring that impossible trajectories would not be considered. Finally, our MCEM algorithm can also be extended by substituting the Markov proposal model with a richer proposal model from which it is straightforward to sample paths conditional on the panel data. Candidates include discrete mixtures of Markov models (Elvira and others, 2019) and phase-type continuous-time Markov chains (Titman and Sharples, 2010). These richer proposal would have the potential of making the importance sampling estimate of the marginal log-likelihood, and thereby the MCEM algorithm as a whole, more robust.
@@ Acknowledgments
The authors would like to thank to Michael P. Fay, Jing Wang, Arthur Winer, and Jennifer Gillman for their invaluable help that greatly improved this manuscript. This work utilized the computational resources of the NIH HPC Biowulf cluster (https://hpc.nih.gov). This project has been funded in whole or in part with federal funds from the National Cancer Institute, National Institutes of Health, under Contract No. 75N91019D00024. The content of this publication does not necessarily reflect the views or policies of the Department of Health and Human Services, nor does mention of trade names, commercial products, or organizations imply endorsement by the U.S. Government. Conflict of Interest: None declared.
@@ A.4.2 MCEM stopping rule
A.4.3 Considerations for MCEM tuning parameters
The computation time of our MCEM algorithm is driven by the optimization steps. The algorithm possesses three tuning parameters specified by the user, which impact the number of optimization steps and the cost of these steps: the tolerance $\epsilon>0$, upper bound limit $10<\gamma<$ 1 and lower bound limit $10<\alpha<$ 1. Decreasing $\epsilon$ and $\gamma$ make the stopping criterion of the MCEM more stringent, resulting in an estimate of the MLE that is closer to the true MLE at the cost of additional optimization steps. Decreasing $\alpha$ increases the frequency with which the target ESS is increased, resulting in the sampling of additional complete paths for the importance sampling estimate of the Q function in the E-step. A larger number of sample paths result in an importance sampling estimate of the Q function that contains more terms and is therefore more expensive to evaluate in the optimization steps. However, a larger number of sample path decreases the uncertainty in the estimate of the Q function, which may result in an earlier termination of the MCEM algorithm.
@@ D.3 Additional results
Appendix E Technical Appendix: Comparison of our proposal distribution for sampling the paths to existing approaches.
E.1 Comparison with Aralis and Brookmeyer (2019)
We compare our sampler for the paths to that in Aralis and Brookmeyer (2019), which we refer to as the A&B sampler. The A&B sampler generates paths from the semi-Markov model via forward simulation and rejects those inconsistent with the observed panel data. Our sampler sample from a surrogate Markov model conditionally on the data. All sample paths are therefore consistent with the data. We then use importance weights to correct for the fact that we sample from a surrogate model.
Table S19 compares the efficiency of these two samplers for fitting a three-state illness-death model to panel data for 500 individuals. Here, efficiency is measured as the ratio of effective sample size to the number of paths sampled: $ESS/M$. A value of $1$ indicates that the sampler is as efficient as Monte Carlo sampler, while a value of, say, $0.1$ indicates that one needs to sample $10$ times more paths to obtain the same amount of information as a Monte Carlo sampler.
Our sampler significantly outperforms the A&B sampler. The efficiency of the A&B sampler decreases as the number of observations per individuals increases, because forward simulation is less likely to generate paths consistent with the observed data, leading to more rejections. In contrast, our approach fares better with richer data, because our sampler conditions on the observed data. Notably, even in the sparser data setting with only 3 observations per individual, our sampler is almost $30$ times more efficient.
The model used for the comparison in Table S19 has only three states, which favors the A&B sampler. The performance of this sampler would decrease for models with more states as it becomes less likely to sample a path consistent with the data. In contrast, the efficiency of our sampler is not impacted by the number of states in the model since it conditions on the observed data.
Table S19: Comparison of the efficiency of our sampler for the paths and that of Aralis and Brookmeyer (2019) (referred to as the A&B sampler) in terms of the ratio of the effective sample size (ESS) over the number of sample paths (M).
Number of observations per individual ESS/M for our sampler ESS/M for the A&B sampler Efficiency gain
$3$ $0.793$ $0.027$ $29.8$
$5$ $0.833$ $0.003$ $308.4$
$10$ $0.923$ $0.0003$ $3077.5$
E.2 Comparison with phase-type Markov models
[…]
Figure S5 compares the fitted models and the data with regard to the prevalence curves in each state. The Markov model, with a constant intensity between the healthy and ill states, yields a prevalence in state 2 that increases too early. To compensate for the early increase, the peak of the prevalence in state 2 is much earlier than the observed peak and the rest of the prevalence is too low after time $t=1$. In contrast, the semi-Markov model with a spline intensity closely matches the observed incidence in state 2. Finally, the results of the phase-type model lies between those of the Markov and spline models. The premature increase in the prevalence in state 2 observed for the Markov and phase-type models is due to the transition intensities for the intensities for transitioning from state 1 and eventually into state 2 being too large near time 0. In contrast, under the spline model, the intensity at $t=0$ can be arbitrarily close to 00.
Table S20: Comparison of the goodness of fit of semi-Markov and phase-type models on simulated data from an illness-death model. The quantity $\ell(\mathbf{Y})-\ell_{M}(\mathbf{Y})$ indicates the absolute gain in log-likelihood over the Markov model with exponential transition intensities (larger is better).
Model $\ell(\mathbf{Y})-\ell_{M}(\mathbf{Y})$ $\Delta$ AIC
Markov 00
Phase-type – 2 latent states $210.9$413.9
Semi-Markov – Deg. 1 splines $\mathbf{247.6}$491.2
Figure S5: Prevalence curves for the illness-death model under a Markov, a phase-type and a semi-Markov multistate model.